# Recognition through Reasoning: Reinforcing Image Geo-localization with Large Vision-Language Models

### v1

@@ Abstract
Previous methods for image geo-localization have typically treated the task as either classification or retrieval, often relying on black-box decisions that lack interpretability. The rise of large vision-language models (LVLMs) has enabled a rethinking of geo-localization as a reasoning-driven task grounded in visual cues. However, two major challenges persist. On the data side, existing reasoning-focused datasets are primarily based on street-view imagery, offering limited scene diversity and constrained viewpoints. On the modeling side, current approaches predominantly rely on supervised fine-tuning, which yields only marginal improvements in reasoning capabilities. To address these challenges, we propose a novel pipeline that constructs a reasoning-oriented geo-localization dataset, MP16-Reason, using diverse social media images. We introduce GLOBE, Group-relative policy optimization for Locatability assessment and Optimized visual-clue reasoning, yielding Bi-objective geo-Enhancement for the VLM in recognition and reasoning. GLOBE incorporates task-specific rewards that jointly enhance locatability assessment, visual clue reasoning, and geolocation accuracy. Both qualitative and quantitative results demonstrate that GLOBE outperforms state-of-the-art open-source LVLMs on geo-localization tasks, particularly in diverse visual scenes, while also generating more insightful and interpretable reasoning trajectories.
@@ 1 Introduction
The Background of Geo-localization. The rapid growth of visual content on social media and mobile devices, has made image geo-localization (determining where an image was taken) increasingly important for downstream applications such as autonomous navigation desouza2002vision and crisis response firmansyah2024empowering . Given that metadata (i.e., GPS coordinates) is frequently unavailable in practice friedland2010cybercasing , predicting geographic location from visual content remains a crucial capability. This demand has led to growing interest in the image geo-localization task hays2008im2gps .
Limitations in Existing Geo-localization Approaches. Traditional image geo-localization approaches fall into two main categories: classification and retrieval. Classification-based methods weyand2016 ; seo2018 ; muller2018 ; pramanick2022 ; clark2023 treat geo-localization as a discrete prediction task, assigning each image to a predefined set of geographical regions or cells. Retrieval-based methods yang2021cross ; zhu2022transgeo ; wang2023fine ; vivanco2023geoclip ; haas2024pigeon ; xia2025fg ; haas2023learning estimate location by comparing the query image to a large geo-tagged reference database, retrieving the closest match in terms of visual features, geographic coordinates, or semantic labels (e.g., city or country names). Although these methods perform well on standard benchmarks, they typically require training on millions of samples and lack interpretability, offering little insight into their underlying reasoning process.
@@
Limitations in LVLM-based Image Geo-localization. Notably, geo-localization requires deeper reasoning than typical vision-language tasks. Success depends on more than recognition, as models must often draw on domain knowledge to infer plausible locations from subtle visual clues such as vegetation, architecture, or language, especially when iconic landmarks are absent. While LVLMs offer a promising path toward such reasoning-driven geo-localization, two fundamental challenges persist, as illustrated in Figure 1. On the data side, existing datasets rarely provide explicit reasoning supervision, such as interpretations of visual evidence and contextual justifications supporting the final location decision. Recent efforts li2024georeasoner ; dou2024gaga ; song2025geolocation to incorporate reasoning into geo-localization datasets have primarily relied on street-view imagery, which offers limited scene diversity and fixed viewpoints. As a result, models trained on such data often struggle to generalize to diverse, real-world visual conditions. On the modeling side, most current approaches depend on supervised fine-tuning with instruction-style data, which tends to encourage pattern replication rather than the development of a grounded understanding of visual-geographic relationships. Without verification mechanisms, these models rely heavily on correlation rather than structured inference, reducing their ability to generalize beyond familiar examples.
@@
How GLOBE Tackles the Challenges. To address these challenges, we propose a novel pipeline for reasoning-aware geo-localization consisting of two main components: (1) constructing a geo-localization dataset from diverse social media images augmented with model-derived reasoning traces, and (2) fine-tuning a vision-language model using Group Relative Policy Optimization (GRPO) for enhanced reasoning. We begin by building MP16-Reason, an extension of MP-16 larson2017benchmarking , which contains user-captured photographs with diverse viewpoints and rich contextual content. To introduce reasoning supervision, we prompt multiple vision-language models Qwen2.5vl ; zhu2025internvl3 ; vivanco2023geoclip to distill the geolocation-related knowledge, including locatability assessments, reasoning trajectories, and predicted locations. To ensure the reliability of these distilled signals, we employ a multi-dimensional verification process that assesses both the alignment between visual evidence and model-generated reasoning, and the consistency across different models through self-verification, thereby filtering out inconsistent or hallucinated outputs. Finally, we fine-tune a pretrained LVLM on the curated dataset using GRPO shao2024deepseekmath , guided by task-specific rewards for locatability, visual grounding, and geolocation accuracy. Our resulting model, GLOBE, achieves state-of-the-art performance among open-source VLMs on geo-localization benchmarks, while producing more interpretable and visually grounded reasoning trajectories, as shown in Figure 2. Our main contributions include:
@@
    Reasoning-Aware Geo-Localization Dataset: We construct MP16-Reason, a diverse geo-localization dataset enriched with image-grounded reasoning supervision that supports model interpretability and generalization.
@@
    GRPO-Based Fine-Tuning: We develop a GRPO-based reinforcement learning framework that fine-tunes LVLMs using task-specific rewards for locatability, visual grounding, and geolocation accuracy, enabling stronger reasoning capabilities compared to traditional supervised fine-tuning.
@@ 2 Related Work
Image Geo-localization. Image geo-localization aims to predict the geographic location of a given image and has broad applications in urban analysis yeh1999urban ; firmansyah2024empowering ; yan2024urbanclip ; ye2019measuring ; ye2019visual , navigation desouza2002vision , and geospatial data mining korting2013geodma ; liu2024unitime ; liang2018geoman ; pan2019urban ; hao2025nature . With advances in multimodal models, research has evolved from classification weyand2016 ; seo2018 ; muller2018 ; pramanick2022 ; clark2023 and retrieval-based methods yang2021cross ; zhu2022transgeo ; wang2023fine ; vivanco2023geoclip ; haas2024pigeon ; xia2025fg ; haas2023learning to generation-based approaches li2024georeasoner ; zhou2024img2loc ; jia2024g3 , which aim to produce location predictions through visual reasoning. Recent studies li2024georeasoner ; zhou2024img2loc ; jia2024g3 have pointed out key limitations of classification (e.g., coarse granularity) and retrieval methods (e.g., dependency on large reference databases), prompting increased interest in generation-based alternatives. Since the introduction of the MediaEval Placing Tasks 2016 (MP-16) dataset by larson2017benchmarking , recent research zhou2024img2loc ; jia2024g3 continues to utilize this dataset to model relationships between visual semantics and geographic locations. In contrast to conventional approaches, current LVLMs liu2023llava ; Qwenvl ; Qwen2vl ; Qwen2.5vl ; chen2024internvl , which are typically pre-trained on large-scale datasets, inherently exhibit significant visual reasoning capabilities. This raises the critical question of whether the continued reliance on millions of labeled samples for supervised fine-tuning remains necessary to effectively adapt these models to specific tasks. In this work, we take a data-centric perspective to investigate how existing large-scale datasets can be leveraged to construct higher-quality training data for fine-tuning LVLMs in the context of image geo-localization.
@@
Large Vision-Language Models. Building upon advancements in LLMs brown2020language ; touvron2023llama ; vicuna2023 ; qwen2023 ; bao2024harnessing ; guo2025deepseek ; he2025skywork , LLaVA liu2023llava was a pioneering approach in this direction, combining a pre-trained vision encoder with an LLM. It demonstrated that joint fine-tuning of visual and textual representations significantly enhances performance in tasks such as image-based question answering lu2022learn ; yue2024mmmu ; lumathvista ; li2024seed . Subsequently, various LVLMs have emerged Qwenvl ; Qwen2vl ; Qwen2.5vl ; chen2024internvl ; achiam2023gpt ; team2023gemini , differing primarily in their visual-language alignment mechanisms and associated architectural trade-offs. Motivated by these recent advancements, our work further investigates the shift of image geo-localization from traditional methods to LVLMs. Specifically, we explore how curated datasets can be effectively leveraged to facilitate more efficient fine-tuning of these models for geo-localization tasks.
@@
We propose a novel pipeline based on the original MP-16 larson2017benchmarking dataset, aiming to advance image geo-localization from single-modal visual recognition to more robust multimodal reasoning. Achieving this objective requires not only powerful models but also well-curated training data that effectively capture geographic clues. Our pipeline for reasoning-aware geo-localization consists of two main components: dataset curation and model fine-tuning. These are implemented in three stages: (1) dataset curation via strong-to-weak distillation & verfication (Section 3.1), (2) reward construction via task-specific supervision (Section 3.2), and (3) model fine-tuning via GRPO-based reinforcement learning (Section 3.3).
@@ 3.1 Dataset Curation: Data Distillation & Verfication
Raw web-scale datasets contain a diverse range of social media images captured from varied perspectives. However, these datasets suffer from substantial noise deng2009imagenet ; li2017webvision ; wei2022learning ; wei2023aggregate ; liu2024automatic , such as close-up shots with limited visual context or generic objects lacking informative localizable clues. To address this issue and select appropriate images for downstream training, we employ multiple vision-language model knowledge distillation for data synthesis and multi-dimensional verification for data curation.
@@
Multiple Vision-Language Models Knowledge Distillation. We utilize multiple vision-language models (e.g., Qwen2.5-VL-72B Qwen2.5vl , InternVL3-78B zhu2025internvl3 , and GeoCLIP vivanco2023geoclip ) to extract locatability judgments, visual clues, and geolocation predictions for each image in the MP-16 larson2017benchmarking dataset, following the approach of li2024georeasoner ; campos2025gaea . As shown in Figure 3, Qwen2.5-VL and InternVL3 produce binary locatability decisions, step-by-step reasoning trajectories, and textual geolocation predictions. GeoCLIP, in contrast, produces latitude-longitude coordinates along with a confidence score that quantifies locatability campos2025gaea . Collectively, these strong models offer complementary signals, which we distill into structured supervision for downstream data curation and reward modeling.
Multi-dimensional Verfication. Following model inference, we perform multi-dimensional verification to curate high-quality data, as illustrated in Figure 3. Initially, we filter out images with negative locatability decisions or low locatability scores. Subsequently, incorrect geolocation predictions are discarded by comparing them against ground-truth annotations. To ensure the reliability of the knowledge distilled from Qwen2.5-VL and InternVL3, we introduce a self-verification step in which the geolocation predictions and reasoning trajectories of both models are compared for each image. Only those samples exhibiting consistent location outputs (e.g., matching city- or country-level predictions) and semantically aligned reasoning chains are retained. This cross-model agreement serves as the reliability proxy in distilled supervision. Furthermore, to enforce visual grounding of the reasoning process, we employ a general-purpose semantic segmentation model cheng2021per to extract both the categories and relative proportions of visual elements within each image. We then assess the consistency between the entities mentioned in the reasoning trajectories and the visual elements identified via segmentation. Through this multi-stage validation pipeline, which combines locatability filtering, self-verification of distilled knowledge, and visual-semantic consistency checks, we curate a robust and trustworthy dataset tailored for downstream tasks.
@@ 3.2 Reward Construction: Task-specific Supervision
Locatability Reward. We develop a binary classification reward model to estimate the localizability of an image based on its visual content. Using the curated dataset $\mathcal{D}$, we train a model $R_{\text{loc}}(I_{i})\in\{0,1\}$ to predict the probability that $y_{i}=1$, i.e., the image is localizable. Accordingly, we define the reward as:
$R_{\text{loc}}(I_{i})=\mathbb{P}(y_{i}=1\mid I_{i};\theta_{\text{loc}}),$ (1)
where $\theta_{\text{loc}}$ denotes the parameters of the reward model. The output score serves both as a reward signal for reinforcement learning and as a soft indicator of an image’s inherent locatability.
Visual Grounding Consistency Reward. To ensure the model-generated reasoning aligns with the actual visual content, we introduce a reward model evaluating entity grounding consistency. For a given sample $(I_{i},r_{i})$ from the curated dataset, let $\hat{r}_{i}$ denote the predicted reasoning. We extract a set of entities $E_{i}=\{e_{1},e_{2},...,e_{n}\}$ from the reasoning trajectory $\hat{r}_{i}$, and a set of visual elements $V_{i}=\{v_{1},v_{2},...,v_{m}\}$ from both the image $I_{i}$ (via semantic segmentation) and the text of $r_{i}$ (via entity extraction). We define a soft matching function $\text{Match}(e_{j},V_{i})\in{0,1}$, which returns 1 if entity $e_{j}$ approximately matches any element in $V_{i}$, allowing for partial lexical or semantic overlap. The visual grounding reward is computed as:
@@
where $\mathbb{I}[\cdot]$ is the indicator function and $\alpha\in[0,1]$ is a weighting factor that controls the importance of city-level correctness, conditional on the country being correct. This reward structure captures the hierarchical nature of geo-tags. A reward of 00 is assigned when the predicted country is incorrect (i.e., $\hat{c}_{i}\neq c_{i}$). If the country is correct but the city is not (i.e., $\hat{c}_{i}=c_{i}$, $\hat{t}_{i}\neq t_{i}$), the model receives a partial reward of $1-\alpha$. A full reward of $1$ is assigned only when both predictions are correct (i.e., $\hat{c}_{i}=c_{i}$, $\hat{t}_{i}=t_{i}$). This tiered design encourages the model to first learn coarse-grained localization before refining its predictions to finer spatial resolutions.

### v2

@@ Abstract
Previous methods for image geo-localization have typically treated the task as either classification or retrieval, often relying on black-box decisions that lack interpretability. The rise of large vision-language models (LVLMs) has enabled a rethinking of geo-localization as a reasoning-driven task grounded in visual cues. However, two major challenges persist. On the data side, existing reasoning-focused datasets are primarily based on street-view imagery, offering limited scene diversity and constrained viewpoints. On the modeling side, current approaches predominantly rely on supervised fine-tuning, which yields only marginal improvements in reasoning capabilities. To address these challenges, we propose a novel pipeline that constructs a reasoning-oriented geo-localization dataset, MP16-Reason, using diverse social media images. We introduce GLOBE, Group-relative policy optimization for Localizability assessment and Optimized visual-cue reasoning, yielding Bi-objective geo-Enhancement for the VLM in recognition and reasoning. GLOBE incorporates task-specific rewards that jointly enhance localizability assessment, visual-cue reasoning, and geolocation accuracy. Both qualitative and quantitative results demonstrate that GLOBE outperforms state-of-the-art open-source LVLMs on geo-localization tasks, particularly in diverse visual scenes, while also generating more insightful and interpretable reasoning trajectories. The data and code are available at https://github.com/lingli1996/GLOBE.
@@ 1 Introduction
The Background of Geo-localization. The rapid growth of visual content on social media and mobile devices, has made image geo-localization (determining where an image was taken) increasingly important for downstream applications such as autonomous navigation desouza2002vision ; mavrogiannis2023core ; nahavandi2025comprehensive and crisis response firmansyah2024empowering . Given that metadata (i.e., GPS coordinates) is frequently unavailable in practice friedland2010cybercasing , predicting geographic location from visual content remains a crucial capability. This demand has led to growing interest in the image geo-localization task hays2008im2gps .
Limitations in Existing Geo-localization Approaches. Traditional image geo-localization approaches fall into two main categories: classification and retrieval. Classification-based methods weyand2016 ; seo2018 ; muller2018 ; pramanick2022 ; clark2023 treat geo-localization as a discrete prediction task, assigning each image to a predefined set of geographical regions or cells. Retrieval-based methods yang2021cross ; zhu2022transgeo ; wang2023fine ; vivanco2023geoclip ; haas2024pigeon ; xia2025fg ; haas2023learning estimate location by comparing the query image to a large geo-tagged reference database, retrieving the closest match in terms of visual features, geographic coordinates, or semantic labels (e.g., city or country names). In practice, retrieval-based approaches can achieve higher accuracy at fine-grained precision jia2024g3 ; jia2025georanker , making them particularly effective when exact localization is required. Although these methods perform well on standard benchmarks, they typically require training on millions of samples and lack interpretability, offering little insight into their underlying reasoning process.
@@
Limitations in LVLM-based Image Geo-localization. Notably, geo-localization requires deeper reasoning than typical vision-language tasks. Success depends on more than recognition, as models must often draw on domain knowledge to infer plausible locations from subtle visual cues such as vegetation, architecture, or language, especially when iconic landmarks are absent. While LVLMs offer a promising path toward such reasoning-driven geo-localization, two fundamental challenges persist, as illustrated in Figure 1. On the data side, existing datasets rarely provide explicit reasoning supervision, such as interpretations of visual evidence and contextual justifications supporting the final location decision. Recent efforts li2024georeasoner ; dou2024gaga ; song2025geolocation to incorporate reasoning into geo-localization datasets have primarily relied on street-view imagery, which offers limited scene diversity and fixed viewpoints. As a result, models trained on such data often struggle to generalize to diverse, real-world visual conditions. On the modeling side, most current approaches depend on supervised fine-tuning with instruction-style data, which tends to encourage pattern replication rather than the development of a grounded understanding of visual-geographic relationships. Without verification mechanisms, these models rely heavily on correlation rather than structured inference, reducing their ability to generalize beyond familiar examples.
@@
How GLOBE Tackles the Challenges. To address these challenges, we propose a novel pipeline for reasoning-driven geo-localization consisting of two main components: (1) constructing a geo-localization dataset from diverse social media images augmented with model-derived reasoning traces, and (2) fine-tuning a vision-language model using Group Relative Policy Optimization (GRPO) for enhanced reasoning. We begin by building MP16-Reason, an extension of MP-16 larson2017benchmarking , which contains user-captured photographs with diverse viewpoints and rich contextual content. To introduce reasoning supervision, we prompt multiple vision-language models Qwen2.5vl ; zhu2025internvl3 ; vivanco2023geoclip to distill the geolocation-related knowledge, including localizability assessments, reasoning trajectories, and predicted locations. To ensure the reliability of these distilled signals, we employ a multi-dimensional verification process that assesses both the alignment between visual evidence and model-generated reasoning, and the consistency across different models through self-verification, thereby filtering out inconsistent or hallucinated outputs. Finally, we fine-tune a pretrained LVLM on the curated dataset using GRPO shao2024deepseekmath , guided by task-specific rewards for localizability, visual grounding, and geolocation accuracy. Our resulting model, GLOBE, achieves state-of-the-art performance among open-source VLMs on geo-localization benchmarks, while producing more interpretable and visually grounded reasoning trajectories, as shown in Figure 2. Our main contributions include:
@@
    Reasoning-Oriented Geo-Localization Dataset: We construct MP16-Reason, a diverse geo-localization dataset enriched with image-grounded reasoning supervision that supports model interpretability and generalization.
@@
    GRPO-Based Fine-Tuning: We develop a GRPO-based reinforcement learning framework that fine-tunes LVLMs using task-specific rewards for localizability, visual grounding, and geolocation accuracy, enabling stronger reasoning capabilities compared to traditional supervised fine-tuning.
@@ 2 Related Work
Image Geo-localization. Image geo-localization aims to predict the geographic location of a given image and has broad applications in urban analysis yeh1999urban ; firmansyah2024empowering ; yan2024urbanclip ; ye2019measuring ; ye2019visual , navigation desouza2002vision , and geospatial data mining korting2013geodma ; liu2024unitime ; liang2018geoman ; pan2019urban ; hao2025nature . General methods like Visual Place Recognition (VPR) arandjelovic2016netvlad ; berton2022rethinking ; lu2024cricavpr ; ali2024boq focus on robustness to challenging variations (e.g., illumination and viewpoint). With advances in multimodal models, research has evolved from classification weyand2016 ; seo2018 ; muller2018 ; pramanick2022 ; clark2023 and retrieval-based methods yang2021cross ; zhu2022transgeo ; wang2023fine ; vivanco2023geoclip ; haas2024pigeon ; xia2025fg ; haas2023learning to generation-based approaches izbicki2019exploiting ; li2024georeasoner ; zhou2024img2loc ; jia2024g3 ; dufour2025around , which aim to produce location predictions through visual reasoning. Recent studies li2024georeasoner ; zhou2024img2loc ; jia2024g3 have pointed out key limitations of classification (e.g., coarse granularity) and retrieval methods (e.g., dependency on large reference databases), prompting increased interest in generation-based alternatives. Since the introduction of the MediaEval Placing Tasks 2016 (MP-16) dataset by larson2017benchmarking , recent research zhou2024img2loc ; jia2024g3 continues to utilize this dataset to model relationships between visual semantics and geographic locations. In contrast to conventional approaches, current LVLMs liu2023llava ; Qwenvl ; Qwen2vl ; Qwen2.5vl ; chen2024internvl , which are typically pre-trained on large-scale datasets, inherently exhibit significant visual reasoning capabilities. This raises the critical question of whether the continued reliance on millions of labeled samples for supervised fine-tuning remains necessary to effectively adapt these models to specific tasks. In this work, we take a data-centric perspective to explore how large-scale datasets can be used to build higher-quality training data for fine-tuning LVLMs in image geo-localization.
@@
Large Vision-Language Models. Building on recent LLM advancements brown2020language ; touvron2023llama ; vicuna2023 ; qwen2023 ; bao2024harnessing ; guo2025deepseek ; he2025skywork , LLaVA liu2023llava demonstrated that combining a vision encoder with an LLM and jointly fine-tuning them improves image-based question answering lu2022learn ; yue2024mmmu ; lumathvista ; li2024seed . Subsequently, various LVLMs have emerged Qwenvl ; Qwen2vl ; Qwen2.5vl ; chen2024internvl ; achiam2023gpt ; team2023gemini , differing primarily in their visual-language alignment mechanisms and associated architectural trade-offs. Motivated by these recent advancements, our work further investigates the shift of image geo-localization from traditional methods to LVLMs. Specifically, we explore how curated datasets can be effectively leveraged to facilitate more efficient fine-tuning of these models for geo-localization tasks.
@@
We propose a novel pipeline based on the original MP-16 larson2017benchmarking dataset, aiming to advance image geo-localization from single-modal visual recognition to more robust multimodal reasoning. Achieving this objective requires not only powerful models but also well-curated training data that effectively capture geographic cues. Our pipeline for reasoning-driven geo-localization consists of two main components: dataset curation and model fine-tuning. These are implemented in three stages: (1) dataset curation via strong-to-weak distillation & verification (Section 3.1), (2) reward construction via task-specific supervision (Section 3.2), and (3) model fine-tuning via GRPO-based reinforcement learning (Section 3.3).
@@ 3.1 Dataset Curation: Data Distillation & Verfication
Raw web-scale datasets contain a diverse range of social media images captured from varied perspectives. However, these datasets suffer from substantial noise deng2009imagenet ; li2017webvision ; wei2022learning ; wei2023aggregate ; liu2024automatic , such as close-up shots with limited visual context or generic objects lacking informative localizable cues. To address this issue and select appropriate images for downstream training, we employ multi-model vision-language knowledge distillation for data synthesis and multi-dimensional verification for data curation.
@@
Multiple Vision-Language Models Knowledge Distillation. We utilize multiple vision-language models (e.g., Qwen2.5-VL-72B Qwen2.5vl , InternVL3-78B zhu2025internvl3 , and GeoCLIP vivanco2023geoclip ) to extract localizability judgments, visual cues, and geolocation predictions for each image in the MP-16 larson2017benchmarking dataset, inspired by izbicki2019exploiting ; li2024georeasoner ; campos2025gaea . The use of three diverse, high-performing VLMs is a deliberate design choice to mitigate model-specific biases. Rather than relying on a single model, which may reflect its own systematic preferences or reasoning patterns, combining multiple VLMs allows us to leverage their consensus and complementarity, thereby enhancing both the robustness and diversity of the distilled signals zhao2024mtms . As shown in Figure 3, Qwen2.5-VL-72B Qwen2.5vl and InternVL3-78B zhu2025internvl3 produce binary localizability decisions, step-by-step reasoning trajectories, and textual geolocation predictions. GeoCLIP vivanco2023geoclip , in contrast, produces latitude-longitude coordinates along with a confidence score that quantifies localizability campos2025gaea . Collectively, these strong models offer complementary signals, which we distill into structured supervision for downstream data curation and reward modeling.
Multi-dimensional Verification. Following model inference, we perform multi-dimensional verification to curate high-quality data, as illustrated in Figure 3. Initially, we filter out images with negative localizability decisions or low localizability scores. Subsequently, incorrect geolocation predictions are discarded by comparing them against ground-truth annotations. To ensure the reliability of the knowledge distilled from Qwen2.5-VL-72B Qwen2.5vl and InternVL3-78B zhu2025internvl3 , we introduce a self-verification step in which the geolocation predictions and reasoning trajectories of both models are compared for each image. Only those samples exhibiting consistent location outputs (e.g., matching city- or country-level predictions) and semantically aligned reasoning chains are retained. This cross-model agreement serves as the reliability proxy in distilled supervision. Furthermore, to enforce visual grounding of the reasoning process, we employ a general-purpose semantic segmentation model cheng2021per to extract both the categories and relative proportions of visual elements within each image. The segmentation produces pixel-level labels across a wide range of semantic categories (e.g., sky, building, road, vegetation, and car), providing a dense understanding of the scene composition. We then assess the consistency between the entities mentioned in the reasoning trajectories and the detected visual elements, ensuring that the reasoning is supported by actual visual evidence rather than coincidental correlations. Through this multi-stage validation pipeline, which combines localizability filtering, self-verification of distilled knowledge, and visual-semantic consistency checks, we curate a robust and trustworthy dataset tailored for downstream tasks.
@@ 3.2 Reward Construction: Task-specific Supervision
Localizability Reward. We introduce a localizability reward to estimate how well an image, together with its predicted reasoning $\hat{r}_{i}$, can support reliable localization. In other words, localizability reflects the joint contribution of the visual content and the reasoning process to the likelihood of correct localization. To this end, we train a LLM-based reward model on the curated dataset $\mathcal{D}$, where the objective is to distinguish whether a given pair $(I_{i},\hat{r}i)$ corresponds to a localizable case ($y_{i}=1$). Instead of using only the image as input, incorporating the predicted reasoning allows the model to exploit semantic cues that indicate the interpretability and consistency of the localization process. Formally, the reward is defined as:
$R_{\text{loc}}(I_{i},\hat{r}_{i})=\mathbb{P}(y_{i}=1\mid I_{i},\hat{r}_{i};\theta_{\text{loc}}),$ (1)
where $\theta_{\text{loc}}$ denotes the parameters of the reward model. The resulting probability score serves as a reward signal for reinforcement learning and as a soft indicator of the localizability of the image–reasoning pair.
Visual Grounding Consistency Reward. To ensure the model-generated reasoning aligns with the actual visual content, we introduce a reward model that evaluates entity grounding consistency. For a given sample $(I_{i},r_{i})$ from the curated dataset, let $\hat{r}_{i}$ denote the predicted reasoning. We extract a set of entities $E_{i}=\{e_{1},e_{2},...,e_{n}\}$ from the reasoning trajectory $\hat{r}_{i}$, and a set of visual elements $V_{i}=\{v_{1},v_{2},...,v_{m}\}$ from both the image $I_{i}$ (via semantic segmentation) and the text of $r_{i}$ (via entity extraction). We define a soft matching function $\text{Match}(e_{j},V_{i})\in{0,1}$, which returns 1 if entity $e_{j}$ approximately matches any element in $V_{i}$, allowing for partial lexical or semantic overlap. The visual grounding reward is computed as:
@@
where $\mathbb{I}[\cdot]$ is the indicator function and $\alpha\in[0,1]$ is a weighting factor that controls the importance of city-level correctness, conditional on the country being correct. This reward structure captures the hierarchical nature of geo-tags. A reward of 0 is assigned when the predicted country is incorrect (i.e., $\hat{c}_{i}\neq c_{i}$). If the country is correct but the city is not (i.e., $\hat{c}_{i}=c_{i}$, $\hat{t}_{i}\neq t_{i}$), the model receives a partial reward of $1-\alpha$. A full reward of $1$ is assigned only when both predictions are correct (i.e., $\hat{c}_{i}=c_{i}$, $\hat{t}_{i}=t_{i}$). This tiered design encourages the model to first learn coarse-grained localization before refining its predictions to finer spatial resolutions.